# Covariant Quantum Error-Correcting Codes with Metrological Entanglement Advantage

### v1

@@ Acknowledgements.
We consider a system with $N$ spin-$s$ degrees of freedom and the ideal probe state $|\psi\rangle=\frac{1}{\sqrt{2}}(|J,M\rangle+|J,-M\rangle)$, where $J=sN$, and the quantum metrology problem where the sensing parameter $\theta$ couples to $\hat{Q}_{z}$, $|\psi_{\theta}\rangle=e^{-i\theta\hat{Q}_{z}}|\psi\rangle$, or $\psi_{\theta}=\frac{1}{2}(|J,M\rangle\langle J,M|+|J,-M\rangle\langle J,-M|+e^{-i2M\theta}|J,M\rangle\langle J,-M|+e^{i2M\theta}|J,-M\rangle\langle J,M|)$. Since $\hat{Q}_{z}|J,M\rangle=M|J,M\rangle$, we see that the QFI of this probe state is $\mathfrak{F}(\psi_{\theta})=4(\langle\hat{Q}_{z}^{2}\rangle-\langle\hat{Q}_{z}\rangle^{2})=4M^{2}$.
Consider the probe state going through an erasure channel $\rho_{\theta}=\mathcal{N}(\psi_{\theta}):=\text{Tr}_{d}[\psi_{\theta}]$, where $\text{Tr}_{d}[...]$ means tracing out $d$ sites. Note that, since the $J=sN$ irrep is totally permutationally symmetric, we can just assume that we trace out the first $d$ sites. To calculate the QFI of $\rho_{\theta}$, we need to find its eigendecomposition. This can be achieved as follows. First, note that we can rewrite
@@
where the matrix $B_{m}$ is represented by the basis $|j_{2},M-m\rangle$ and $|j_{2},-M-m\rangle$. We also see that $\rho_{\theta}$ can be interpreted as an ensemble labeled by $m=-j_{1},\cdots,j_{1}$, where the $\theta$ information is encoded in the coherence between the states $|j_{2},M\!-\!m\rangle$ and $|j_{2},-M\!-\!m\rangle$, which have a magnetic quantum number separation $2M$. The eigenvalues of the $2\times 2$ matrix $B_{m}$ are $\frac{1}{2}(a_{m}^{2}+b_{m}^{2})$ and 00, and the corresponding eigenvectors are
@@
We propose a simple local measurement for the ideal probe state $|\psi\rangle=\frac{1}{\sqrt{2}}(|J,M\rangle+|J,-M\rangle)$ which saturates the quantum Cramér-Rao bound. Specifically, we consider the observable $\mathbf{D}:=\bigotimes_{j=1}^{N}D_{j}$, where $D_{j}$ satisfies $D_{j}\hat{q}^{z}_{j}D_{j}=-\hat{q}^{z}_{j}$. This is a local measurement since one can measure $\mathbf{D}$ by measuring $D_{j}$ locally and multiplying the outcomes. For example, in the basis $\hat{q}^{z}_{j}|\sigma\rangle=\sigma|\sigma\rangle$, where $\sigma=-s,\cdots s$, one can pick $D_{j}=\sum_{\sigma=-s}^{s}|s\rangle\langle-s|$, which would give $\mathbf{D}^{2}=I$. The unknown parameter $\theta$ can therefore be obtained by measuring $\langle\mathbf{D}\rangle=\langle\psi_{\theta}|\mathbf{D}|\psi_{\theta}\rangle=\cos(2M\theta)$, and we have $\theta=\arccos(\langle\mathbf{D}\rangle)/(2M)$. The variance is $\sigma_{\mathbf{D}}^{2}=\langle\mathbf{D}^{2}\rangle-\langle\mathbf{D}\rangle^{2}=1-\cos^{2}(2M\theta)$, which gives the standard error of the measurement as $\Delta\mathbf{D}=\sigma_{\mathbf{D}}/\sqrt{\nu}$, where $\nu$ is the number of samples. So we have $\Delta\theta=\frac{\Delta\mathbf{D}}{2M}\frac{\arccos(x)}{dx}|_{x=\langle\mathbf{D}\rangle}=\frac{1}{2M\sqrt{\nu}}$, which indeed saturates the quantum Cramér-Rao bound since QFI $\mathfrak{F}(\psi_{\theta})=4M^{2}$.
[…]
$\Delta\theta=\frac{\sigma_{\mathbf{D}^{\prime}}}{\sqrt{\nu}2MA\sin(2M\theta)}\sim\frac{1}{2MA}~{}.$ (S51)
@@
and the asymptotic expansion for $A$ can be carried out like in App. S5. We expect $A=1-O(j_{1}M^{2}/J^{2})$, which gives us $\Delta\theta\sim M^{-1}=N^{-b}$. This scaling matches with the QFI scaling in the red-shaded parameter regime of Fig. 1(b) in the main text.
A global measurement is generally hard to implement so it is usually more desirable to use a local measurement. Below, we show that the local measurement $\mathbf{D}_{\bar{d}}:=\bigotimes_{j=d+1}^{N}D_{j}$ can be used to obtain $\theta$ with a $\Delta\theta$ scaling beyond the standard quantum limit, although it does not saturate the scaling given by the QFI. Note that, for $\rho_{\theta}=\sum_{m=-j_{1}}^{j_{1}}\lambda_{m}|m\rangle\langle m|$, if $j_{1}$ is a half-integer, then we can throw out one more qudit so that $j_{1}$ is an integer. We thus see that $\mathbf{D}_{\bar{d}}$ will pick up the signal from the coherence between states $|j_{2}\!,\!M\!-\!m\rangle$ and $|j_{2}\!,\!-M-m\rangle$ only when $m=0$. More explicitly, we have $\langle\mathbf{D}_{\bar{d}}\rangle=\text{Tr}[\rho_{\theta}\mathbf{D}_{\bar{d}}]=a_{0}b_{0}\cos(2M\theta)$. Comparing this result with Eq. (S51) and noting that $a_{0}b_{0}\sim(j_{1})^{-\frac{1}{2}}$, we obtain $\Delta\theta\sim\sqrt{j_{1}}/M\sim N^{-(b-c/2)}$ if $j_{1}\sim N^{c}$ and $M\sim N^{b}$, exhibiting a scaling beyond the standard quantum limit if $12b-c>$ 1, together with the conditions $b>c$ and $c<2(1-b)$. For $c=0$, the local measurement would give a variance scaling matching the QFI scaling.
S8 Spin-1 XY-Dzyaloshinskii-Moriya Scar States as AQECC
In this appendix, we show that the quantum many-body scar states in the Spin-1 XY-Dzyaloshinskii-Moriya model [65, 66, 67] can form an AQECC. While the original model considers a system with $N$ spin-$1$ degrees of freedom, here we generalize it to $N$ spin-$s$ degrees of freedom. Specifically, we consider the spin-$s$ local Hilbert space with the basis $\{|m\rangle,m=-s,\cdots,s\}$. We define the operators
@@
where $\phi_{j}$ is some site-dependent phase. It is easy to verify that these operators generate the $SU(2)$ algebra $[J^{+},J^{-}]=2J^{z}$ and $[J^{z},J^{\pm}]=\pm J^{\pm}$. The quantum many-body scar states are $|S_{M}\rangle:=Z_{M}(J^{+})^{M}|S_{0}\rangle$, where $|S_{0}\rangle=\bigotimes_{j=1}^{N}|-\!s\rangle_{j}$ and $Z_{M}$ is the normalization constant. The scar states in the Spin-1 XY-Dzyaloshinskii-Moriya model correspond to $s=1$ with some specific choice of $\phi_{j}$.
In fact, these scar states are Dicke states in disguise. This can be seen by redefining $|s\rangle_{j}\rightarrow e^{-\phi_{j}/2}|s\rangle_{j}$ and $|-s\rangle_{j}\rightarrow e^{\phi_{j}/2}|s\rangle_{j}$, which can be achieved by applying a tensor product of single-site unitaries $U=\bigotimes_{j=1}^{N}u_{j}$, where $u_{j}=|s\rangle\langle s|_{j}e^{-\phi_{j}/2}+|\!-\!s\rangle\langle\!-\!s|_{j}e^{\phi_{j}/2}+\sum_{m=-s+1\cdots s-1}|m\rangle\langle m|_{j}$. We therefore see that $|S_{M}\rangle$ are Dicke states $|J\!,\!M\rangle$ ($s=\frac{1}{2}$, $J=\frac{N}{2}$) embedded in a larger local Hilbert space, up to a tensor product of single-site unitary. Therefore, we conclude that the quantum many-body scar states $\mathfrak{C}=\text{span}\{|S_{M}\rangle;M=M_{\mathrm{min}},M_{\mathrm{min}}+\Delta,\cdots,M_{\mathrm{max}}\}$ forms an AQECC against $d$-local noise or erasures at known locations if we take $\Delta\geq 2sd+1$. The code parameters and the inaccuracy bound against different errors can be readily obtained from Theorems 1 and 2 with $s=\frac{1}{2}$, $J=\frac{N}{2}$.

### v2

@@ Acknowledgements.
We consider a system with NN spin-$s$ degrees of freedom and the ideal probe state $|\psi\rangle=\frac{1}{\sqrt{2}}(|J,M\rangle+|J,-M\rangle)$, where $J=sN$, and the quantum metrology problem where the sensing parameter $\theta$ couples to $\hat{Q}_{z}$, $|\psi_{\theta}\rangle=e^{-i\theta\hat{Q}_{z}}|\psi\rangle$, or $\psi_{\theta}=\frac{1}{2}(|J,M\rangle\langle J,M|+|J,-M\rangle\langle J,-M|+e^{-i2M\theta}|J,M\rangle\langle J,-M|+e^{i2M\theta}|J,-M\rangle\langle J,M|)$. Since $\hat{Q}_{z}|J,M\rangle=M|J,M\rangle$, we see that the QFI of this probe state is $\mathfrak{F}(\psi_{\theta})=4(\langle\hat{Q}_{z}^{2}\rangle-\langle\hat{Q}_{z}\rangle^{2})=4M^{2}$.
Consider the probe state going through an erasure channel $\rho_{\theta}=\mathcal{N}(\psi_{\theta}):=\mathrm{Tr}_{d}[\psi_{\theta}]$, where $\mathrm{Tr}_{d}[$...] means tracing out $d$ sites. Note that, since the $J=sN$ irrep is totally permutationally symmetric, we can just assume that we trace out the first $d$ sites. To calculate the QFI of $\rho_{\theta}$, we need to find its eigendecomposition. This can be achieved as follows. First, note that we can rewrite
@@
where the matrix $B_{m}$ is represented by the basis $|j_{2},M-m\rangle$ and $|j_{2},-M-m\rangle$. We also see that $\rho_{\theta}$ can be interpreted as an ensemble labeled by $m=-j_{1},\cdots,j_{1}$, where the $\theta$ information is encoded in the coherence between the states $|j_{2},M\!-\!m\rangle$ and $|j_{2},-M\!-\!m\rangle$, which have a magnetic quantum number separation $2M$. The eigenvalues of the $2\times 2$ matrix $B_{m}$ are $\frac{1}{2}(a_{m}^{2}+b_{m}^{2})$ and 0, and the corresponding eigenvectors are
@@
We propose a simple local measurement for the ideal probe state $|\psi\rangle=\frac{1}{\sqrt{2}}(|J,M\rangle+|J,-M\rangle)$ which saturates the quantum Cramér-Rao bound. Specifically, we consider the observable $\mathbf{D}:=\bigotimes_{j=1}^{N}D_{j}$, where $D_{j}$ satisfies $D_{j}\hat{q}^{z}_{j}D_{j}=-\hat{q}^{z}_{j}$. This is a local measurement since one can measure $\mathbf{D}$ by measuring $D_{j}$ locally and multiplying the outcomes. For example, in the basis $\hat{q}^{z}_{j}|\Gamma\rangle=\Gamma|\Gamma\rangle$, where $\Gamma=-s,\cdots s$, one can pick $D_{j}=\sum_{\Gamma=-s}^{s}|s\rangle\langle-s|$, which would give $\mathbf{D}^{2}=I$. The unknown parameter $\theta$ can therefore be obtained by measuring $\langle\mathbf{D}\rangle=\langle\psi_{\theta}|\mathbf{D}|\psi_{\theta}\rangle=\cos(2M\theta)$, and we have $\theta=\arccos(\langle\mathbf{D}\rangle)/(2M)$. The variance is $\Gamma_{\mathbf{D}}^{2}=\langle\mathbf{D}^{2}\rangle-\langle\mathbf{D}\rangle^{2}=1-\cos^{2}(2M\theta)$, which gives the standard error of the measurement as $\Delta\mathbf{D}=\Gamma_{\mathbf{D}}/\sqrt{\nu}$, where $\nu$ is the number of samples. So we have $\Delta\theta=\frac{\Delta\mathbf{D}}{2M}\frac{\arccos(x)}{dx}|_{x=\langle\mathbf{D}\rangle}=\frac{1}{2M\sqrt{\nu}}$, which indeed saturates the quantum Cramér-Rao bound since QFI $\mathfrak{F}(\psi_{\theta})=4M^{2}$.
For the partially-erased probe state $\rho_{\theta}=\sum_{m=-j_{1}}^{j_{1}}\lambda_{m}|m\rangle\langle m|$, recall that the $\theta$ information is encoded in the coherence between the states $|j_{2},M\!-\!m\rangle$ and $|j_{2},-M\!-\!m\rangle$. A global measurement such as $\mathbf{D}^{\prime}=\sum_{m=-j_{1}}^{j_{1}}|j_{2},M\!-\!m\rangle\langle j_{2},-M\!-\!m|+\text{h.c.}$ would be a candidate to estimate the unknown $\theta$. Indeed, we see $\langle\mathbf{D}^{\prime}\rangle=\mathrm{Tr}[\rho_{\theta}\mathbf{D}^{\prime}]=\sum_{m=-j_{1}}^{j_{1}}a_{m}b_{m}\cos(2M\theta)=A\cos(2M\theta)$, where $A:=\sum_{m=-j_{1}}^{j_{1}}a_{m}b_{m}$. Since $\mathbf{D}^{\prime 2}=\sum_{m=-j_{1}}^{j_{1}}I_{m}$, we have $\Gamma_{\mathbf{D}^{\prime}}^{2}=1-A^{2}\cos^{2}(2M\theta)$, and
$\Delta\theta=\frac{\Gamma_{\mathbf{D}^{\prime}}}{\sqrt{\nu}2MA\sin(2M\theta)}\sim\frac{1}{2MA}~.$ (S73)
@@
and the asymptotic expansion for $A$ can be carried out like in Sec. S6. We expect $A=1-O(j_{1}M^{2}/J^{2})$, which gives us $\Delta\theta\sim M^{-1}=N^{-b}$. This scaling matches with the QFI scaling in the red-shaded parameter regime of Fig. 1(b) in the main text.
A global measurement is generally hard to implement so it is usually more desirable to use a local measurement. Below, we show that the local measurement $\mathbf{D}_{\bar{d}}:=\bigotimes_{j=d+1}^{N}D_{j}$ can be used to obtain $\theta$ with a $\Delta\theta$ scaling beyond the standard quantum limit, although it does not saturate the scaling given by the QFI. Note that, for $\rho_{\theta}=\sum_{m=-j_{1}}^{j_{1}}\lambda_{m}|m\rangle\langle m|$, if $j_{1}$ is a half-integer, then we can throw out one more qudit so that $j_{1}$ is an integer. We thus see that $\mathbf{D}_{\bar{d}}$ will pick up the signal from the coherence between states $|j_{2}\!,\!M\!-\!m\rangle$ and $|j_{2}\!,\!-M-m\rangle$ only when $m=0$. More explicitly, we have $\langle\mathbf{D}_{\bar{d}}\rangle=\mathrm{Tr}[\rho_{\theta}\mathbf{D}_{\bar{d}}]=a_{0}b_{0}\cos(2M\theta)$. Comparing this result with Eq. (S73) and noting that $a_{0}b_{0}\sim(j_{1})^{-\frac{1}{2}}$, we obtain $\Delta\theta\sim\sqrt{j_{1}}/M\sim N^{-(b-c/2)}$ if $j_{1}\sim N^{c}$ and $M\sim N^{b}$, exhibiting a scaling beyond the standard quantum limit if 2bc>12b-c>1, together with the conditions $b>c$ and c<2(1b)c<2(1-b). For $c=0$, the local measurement would give a variance scaling matching the QFI scaling.
S9 A Recovery Channel for AQECC
While a recovery operation is not required for metrological applications, it is nevertheless insightful to understand how a recovery channel can be constructed for our AQECC in the context of quantum error correction. Refs. [18, 71] provide a general framework for constructing nearly optimal recovery channels, which we briefly outline below.
First, we calculate $\lambda_{ij}$ in the Knill-Laflamme condition. While in our earlier analysis, we used $\lambda_{ij}=\langle J,M_{\mathrm{min}}|K_{i}^{\dagger}{K_{j}}|J,M_{\mathrm{min}}\rangle$, in fact, any $M$ satisfying $M_{\mathrm{min}}\leq M\leq M_{\mathrm{max}}$ yields an equally good choice for $\lambda_{ij}=\langle J,M|K_{i}^{\dagger}{K_{j}}|J,M\rangle$ for the purpose of constructing a recovery channel. One then considers the following minimization problem
$F_{\rho}=\min_{\rho}\mathrm{Tr}\sqrt{\Phi_{\rho}(\mathbf{I})}~,$ (S75)
where
$\Phi_{\rho}(\mathbf{I})=\sum_{ij}K_{i}\rho^{2}K_{j}^{\dagger}\lambda_{ij}~,$ (S76)
and the minimization is over the density matrix $\rho=\sum_{a,b=1}^{2^{k}}\rho_{ab}|J\!,\!M_{a}\rangle\langle J\!,\!M_{b}|$ in the logical space.
Suppose $\rho_{0}$ is the density matrix which minimizes the above equation. If $\rho_{0}$ is unique and of full rank, then the following recovery channel is nearly optimal:
$\mathcal{R}(\tau)=\Phi_{\rho_{0}}^{\dagger}[\Phi_{\rho_{0}}(\mathbf{I})^{-\frac{1}{2}}\tau\Phi_{\rho_{0}}(\mathbf{I})^{-\frac{1}{2}}]+\mathcal{T}(\tau)~,$ (S77)
where
$\Phi_{\rho}^{\dagger}(\bullet)=\sum_{ij}\lambda_{ij}[\rho E_{j}^{\dagger}(\bullet)E_{i}\rho]~,$ (S78)
and $\mathcal{T}(\tau)$ is some completely positive map chosen to ensure that $\mathcal{R}(\tau)$ is trace-preserving.
S10 Spin-1 XY-Dzyaloshinskii-Moriya Scar States as AQECC
In this appendix, we show that the quantum many-body scar states in the spin-1 XY-Dzyaloshinskii-Moriya model [72, 73, 74] can form an AQECC. While the original model considers a system with NN spin-11 degrees of freedom, here we generalize it to NN spin-$s$ degrees of freedom. Specifically, we consider the spin-$s$ local Hilbert space with the basis $\{|m\rangle,m=-s,\cdots,s\}$. We define the operators
@@
where $\phi_{j}$ is some site-dependent phase. It is easy to verify that these operators generate the $SU(2)$ algebra $J^{+},J^{-}]=2J^{z}$ and $J^{z},J^{\pm}]=\pm J^{\pm}$. The quantum many-body scar states are $|S_{M}\rangle:=Z_{M}(J^{+})^{M}|S_{0}\rangle$, where $|S_{0}\rangle=\bigotimes_{j=1}^{N}|-\!s\rangle_{j}$ and $Z_{M}$ is the normalization constant. The scar states in the spin-1 XY-Dzyaloshinskii-Moriya model correspond to $s=1$ with some specific choice of $\phi_{j}$.
In fact, these scar states are Dicke states in disguise. This can be seen by redefining $|s\rangle_{j}\rightarrow e^{-\phi_{j}/2}|s\rangle_{j}$ and $|-s\rangle_{j}\rightarrow e^{\phi_{j}/2}|s\rangle_{j}$, which can be achieved by applying a tensor product of single-site unitaries $U=\bigotimes_{j=1}^{N}u_{j}$, where $u_{j}=|s\rangle\langle s|_{j}e^{-\phi_{j}/2}+|\!-\!s\rangle\langle\!-\!s|_{j}e^{\phi_{j}/2}+\sum_{m=-s+1\cdots s-1}|m\rangle\langle m|_{j}$. We therefore see that $|S_{M}\rangle$ are Dicke states $|J\!,\!M\rangle$ ($s=\frac{1}{2}$, $J=\frac{N}{2}$) embedded in a larger local Hilbert space, up to a tensor product of single-site unitary. Therefore, we conclude that the quantum many-body scar states $\mathfrak{C}=\text{span}\{|S_{M}\rangle;M=M_{\mathrm{min}},M_{\mathrm{min}}+\Delta,\cdots,M_{\mathrm{max}}\}$ form an AQECC against $d$-local noise or erasures at known locations if we take $\Delta\geq 2sd+1$. The code parameters and the inaccuracy bound against different errors can be readily obtained from Theorems 1 and 2 with $s=\frac{1}{2}$, $J=\frac{N}{2}$.